# LayerAct: Advanced Activation Mechanism for Robust Inference of CNNs

### v4

@@ Abstract
Codehttps://github.com/KihyukYoon/LayerAct
Appendixhttps://github.com/KihyukYoon/LayerAct
@@ Introduction
Most existing activation functions follow element-level activation (ElementAct) mechanisms, meaning the functions apply independently to each element of the input. However, our analysis identifies a limitation in ElementAct functions: their robustness varies across samples. This is because their robustness relies on the saturation state, where the activation output for a specific range converges to a certain value. For example, in a layer with Rectified Linear Unit (ReLU; Nair and Hinton (2010)), the elements of the output remain unaffected by input shifts only when they are in the saturation state, specifically when the input and shifted input of the elements are smaller than zero. Consequently, existing activation functions can ensure robustness only when a sufficient number of elements are in the saturation state, not when there are fewer elements in that state.
@@
Experimental analysis with the MNIST dataset demonstrates that LayerAct functions have the following properties: (1) the mean activation of LayerAct functions is zero-like, and (2) the output fluctuation due to noisy input is smaller with these functions than with ElementAct functions. We compared the performance of the residual networks (ResNets; He et al. (2016)) with LayerAct functions to those with other ElementAct functions on three image classification tasks. The results on noisy datasets demonstrate that LayerAct functions are superior to other ElementAct functions. Furthermore, ResNet50 with LayerAct functions also showed superior performance on both clean and noisy ImageNet datasets compared to other functions.
@@
where $s$ is increasing and $s\left(y_{i}\right)>0$ if $y_{i}>0$. For example, the activation scale function for the $i^{th}$ element in the sigmoid weighted linear unit (SiLU; Elfwing, Uchibe, and Doya (2018)) is presented as follows:
@@
Table 1 presents the classification performance of the ResNet20 variants with different normalization on both CIFAR10 and CIFAR10-C (Hendrycks and Dietterich 2019). The networks were trained on CIFAR10. The CIFAR10-C dataset serves as a noise-robustness benchmark, incorporating 19 out-of-distribution (OOD) corruptions of CIFAR10. This dataset includes a total of 19 distinct corruptions, each with five levels of severity, organized into five categories: noise, blur, digital, weather, and extra. We evaluated the models using two activation functions: ReLU and SiLU. The table reports the average mean accuracy over 30 runs, with ”Norm” and ”Act” denoting the normalization and activation, respectively. The experimental results show that the use of LN alone leads to inferior performance on the clean dataset, CIFAR10. Specifically, networks with LN alone showed inferior results compared to those without normalization, which aligns with findings from previous studies.
LN and other batch-free normalizations have been shown to block the negative effect of BN on robustness of networks. Interestingly, our experiments reveal a detailed robustness advantage of LN, which can lead to robust activation, and has not been emphasized in previous studies. On the noisy dataset, CIFAR10-C, the networks incorporating LN demonstrated better performance than those without LN. Notably, when LN is applied immediately before the activation layers, as in ResNet20-LN and ResNet20-(BN, LN), these networks outperformed their counterparts, ResNet20-None and ResNet20-(LN, BN), on noisy datasets, despite underperforming on the clean dataset. This suggests that LN may provide a robustness benefit, which is closely related to activation.
@@
For the stability of learning and inference, ensuring the continuity of activation outputs across the entire space is essential. While traditional ElementAct functions, such as ReLU, leaky ReLU (LReLU; Maas et al. (2013)), and PReLU (He et al. 2015), do not require a continuous activation scale at zero (given that the activation output $y_{i}s\left(y_{i}\right)$ remains continuous at zero), LayerAct functions need special consideration, as the activation output $y_{i}s\left(n_{i}\right)$ is discontinuous if the activation scale function is not continuous.
@@
In previous sections, we discussed that LN tends to produce homogeneous activation outputs, as the mean and variance of $n_{i}$ are similar across all samples, which leads to inferior performance. In contrast, LayerAct functions are designed to preserve these diverse representations by transferring the statistics of the activation input $y$ to the activation output $a$.
This difference between activations with LN and LayerAct stems from the difference in the activation output, $n_{i}s\left(n_{i}\right)$ and $y_{i}s\left(n_{i}\right)$, respectively. This difference in approaches demonstrates LayerAct’s ability to maintain the input statistics within the activation outputs, resulting in much more diverse representations of activation compared to activation with LN. For details, see Appendices C and D.
@@
In this section, we present the experimental analysis and classification performance of LayerAct. First, we verify the important properties of LayerAct with the MNIST dataset. Next, we primarily evaluate the classification performance of LayerAct functions on three image benchmark datasets for both clean and noisy cases, and one clean medical image dataset. For the detailed experimental setting to ensure reproducibility, see Appendix F.
@@
Figure 3: Plots of the last layer representation in ResNet20 with LA-SiLU, utilizing no normalization (left) and BN (right).
Figure 4: Distribution of activation output fluctuation due to noise with different noise distribution.
@@
Figure 4 shows the distribution of the activation fluctuation with different noise distributions. Although the fluctuation distribution of the activation input was similar (see Appendix K), LayerAct functions have a significantly smaller mean and variance of activation fluctuation among the samples than any other ElementAct function in all cases. The decrease in variance is remarkable, showing that LayerAct functions are robust for all samples. Moreover, the ElementAct functions that ensure a zero-like activation mean with one-sided saturation such as SiLU or HardSiLU showed slightly larger activation fluctuations than those of ReLU or LReLU when the noise had a large mean. However, LayerAct functions maintained lower fluctuations in both cases.
Figure 5: Distribution of the activation output means of the elements in a trained network on MNIST at epochs $1$ and $40$.
@@
Figure 5 shows the distribution of the activation output means of the single-layer network trained on the MNIST dataset at $1$ and $40$ epochs. The distributions did not change after $40$ epochs. LayerAct functions maintain a zero-like activation mean for all epochs. Our experimental results indicate that LayerAct allows similar (before epoch $20$) or larger (after epoch $40$) negative outputs compared to other ElementAct functions. Thus, LA-SiLU and LA-HardSiLU can achieve a more zero-like activation mean than other activation functions.
@@
We demonstrate the classification performance of LayerAct functions on four image datasets: CIFAR10, CIFAR100, ImageNet (Russakovsky et al. 2015), and a medical image dataset. We used ResNet (He et al. 2016) as the network for our experiments on the three benchmark datasets. In our experiments on CIFAR10, CIFAR100, and ImageNet, we utilized BN. We compared LayerAct functions with ReLU, LReLU, PReLU, SiLU, HardSiLU, Mish (Misra 2020), GELU (Hendrycks and Gimpel 2023), and ELU (Clevert, Unterthiner, and Hochreiter 2016).
The tables report the mean accuracy over 30 runs except for the experiments on ImageNet and the medical image dataset; the best results are underlined and bolded, while the second best are bolded. In the main manuscript, we present the experimental results of ResNet20 on CIFAR10 and ResNet50 and ResNet101 on ImageNet. For more experimental results, see Appendix J.
@@
The column named clean of Table 2 presents the classification performance of ResNet20 with both LayerAct functions and ElementAct functions, benchmarked on the clean CIFAR10. LA-SiLU achieved the best performance among the activation functions. In other experiments, with ResNet32 and ResNet44 on clean CIFAR10 (Appendix J), networks with GELU achieved better results in two cases: ResNet32 on CIFAR10 and ResNet44 on CIFAR10. In the remaining combinations of networks and datasets, LA-SiLU outperformed in a significant majority of cases, especially in 39 out of 48 cases according to the $p$-value from a T-test or a Wilcoxon signed-rank test, indicating statistical significance.
@@
Table 3 shows the performance of ResNet50 and ResNet101 with LayerAct functions and other ElementAct functions for comparison with clean and noisy ImageNet datasets. For the column named ”clean,” we report the accuracy of 10-crop testing on the validation dataset. We used the out-of-distribution benchmark dataset, ImageNet-C (Hendrycks and Dietterich 2019), which has the same corruptions as CIFAR-C datasets. The networks with LayerAct functions outperformed those with other activation functions on all datasets. The ResNet50 with LayerAct functions, even with LA-HardSiLU which showed worse performance on the clean CIFAR10 and CIFAR100 datasets compared to SiLU or LReLU, outperformed other activation functions on clean ImageNet. Based on these results, we trained deeper networks, ResNet101, utilizing ReLU and LA-SiLU as activations. LA-SiLU maintained its superior robustness.

### v5

@@ Abstract
Code & Appendixhttps://github.com/KihyukYoon/LayerAct
@@ Introduction
Most existing activation functions follow element-level activation (ElementAct) mechanisms, meaning the functions apply independently to each element of the input. However, we identified a limitation in the method: their robustness varies across samples. This is because their robustness relies on the saturation state, where the activation output for a specific range converges to a certain value. For example, in a layer with ReLU (Nair and Hinton 2010), the elements of the output remain unaffected by input shifts only when they are in the saturation state, specifically when the input and shifted input of the elements are smaller than zero. Consequently, existing activation functions can ensure robustness only when a sufficient number of elements are in the saturation state, not when there are fewer elements in that state.
@@
Experimental analysis with the MNIST dataset demonstrates that LayerAct functions have the following properties: (1) the mean activation of LayerAct functions is zero-like, and (2) the output fluctuation due to noisy input is smaller with these functions than with ElementAct functions. We compared the performance of the residual networks (ResNets; He et al. (2016)) with LayerAct functions to those with other ElementAct functions on three image classification tasks.
@@
where $s$ is increasing and $s\left(y_{i}\right)>0$ if $y_{i}>0$. For example, the activation scale function for the $i^{th}$ element in the SiLU (Elfwing, Uchibe, and Doya 2018) is presented as follows:
@@
Table 1 demonstrates the performance of the ResNet20 variants with different normalization on both CIFAR10 and CIFAR10-C (Hendrycks and Dietterich 2019). The networks were trained on CIFAR10. The CIFAR10-C dataset serves as a noise-robustness benchmark, incorporating 19 out-of-distribution (OOD) corruptions of CIFAR10. This dataset includes a total of 19 distinct corruptions, each with five levels of severity, organized into five categories: noise, blur, digital, weather, and extra. We evaluated the models using two activation functions: ReLU and SiLU. The table reports the average mean accuracy over 30 runs, with ”Norm” and ”Act” denoting the normalization and activation, respectively. The experimental results show that the use of LN alone leads to inferior performance on the clean dataset, CIFAR10. Specifically, networks with LN alone showed inferior results compared to those without normalization, which aligns with findings from previous studies.
LN and other batch-free normalizations have been shown to block the negative effect of BN on robustness of networks (huang2022delving). Interestingly, our experiments reveal a detailed robustness advantage of LN, which can lead to robust activation, and this has not been emphasized in previous studies. On the noisy dataset, CIFAR10-C, the networks incorporating LN demonstrated better performance than those without LN. Notably, when LN is applied immediately before the activation layers, as in ResNet20-LN and ResNet20-(BN, LN), these networks outperformed their counterparts, ResNet20-None and ResNet20-(LN, BN), on noisy datasets, despite underperforming on the clean dataset. This suggests that LN may provide a robustness benefit, which is closely related to activation.
@@
For the stability of learning and inference, ensuring the continuity of activation outputs across the entire space is essential. While traditional ElementAct functions, such as ReLU, LReLU (Maas et al. 2013), and PReLU (He et al. 2015), do not require a continuous activation scale at zero (given that the activation output $y_{i}s\left(y_{i}\right)$ remains continuous at zero), LayerAct functions need special consideration, as the activation output $y_{i}s\left(n_{i}\right)$ is discontinuous if the activation scale function is not continuous.
@@
We discussed that LN tends to produce homogeneous activation outputs, as the mean and variance of $n_{i}$ are similar across all samples. In contrast, LayerAct functions are designed to preserve these diverse representations by transferring the statistics of the activation input $y$ to the activation output $a$. This difference is from the different activation outputs of activation with LN and LayerAct, $n_{i}s\left(n_{i}\right)$ and $y_{i}s\left(n_{i}\right)$, respectively. Due its unique output, LayerAct can maintain the input statistics within the activation outputs, resulting in much more diverse representations. For details, see Appendices C and D.
@@
In this section, we present the experimental analysis and results of LayerAct. For the detailed experimental setting to ensure reproducibility, see Appendix F.
@@
Figure 3: Plots of the last layer representation in ResNet20 with LA-SiLU, with no normalization (left) and BN (right).
Figure 4: Distribution of activation output fluctuation.
@@
Figure 4 shows the distribution of the activation fluctuation with a noise distribution. Although the fluctuation distribution of the activation input was similar (see Appendix K), LayerAct functions have a significantly smaller mean and variance of activation fluctuation among the samples than any other ElementAct function in all cases. The decrease in variance is remarkable, showing that LayerAct functions are robust for all samples. Moreover, the ElementAct functions that ensure a zero-like activation mean with one-sided saturation such as SiLU or HardSiLU showed slightly larger activation fluctuations than those of ReLU or LReLU when the noise had a large mean. However, LayerAct functions maintained lower fluctuations in both cases.
Figure 5: Distribution of the activation output means of the elements in a trained network on MNIST at epoch $40$.
@@
Figure 5 demonstrates the distribution of the activation output means of the single-layer network trained on the MNIST dataset at the epoch $40$. The distributions did not change after the epoch $40$. LayerAct functions maintain a zero-like activation mean for all epochs. Our experimental results indicate that LayerAct allows similar (before epoch $20$) or larger (after epoch $40$) negative outputs compared to other ElementAct functions. Thus, LA-SiLU and LA-HardSiLU can achieve a more zero-like activation mean than other activation functions.
@@
We demonstrate the classification performance of LayerAct on CIFAR10, ImageNet (Russakovsky et al. 2015), and a medical image dataset. We used ResNet (He et al. 2016) as the network for our experiments on the three benchmark datasets. In our experiments on CIFAR10 and ImageNet, we utilized BN. We compared LayerAct with ReLU, LReLU, PReLU, SiLU, HardSiLU, Mish (Misra 2020), GELU (Hendrycks and Gimpel 2023), and ELU (Clevert, Unterthiner, and Hochreiter 2016). In the tables, The best results are underlined and bolded, and the second best are bolded. In the main manuscript, we present the results of ResNet20 on CIFAR10 and ResNet50 and ResNet101 on ImageNet. For more experimental results, including on CIFAR100, see Appendix J.
@@
The column named clean of Table 2 presents the classification performance of ResNet20 with both LayerAct functions and ElementAct functions, benchmarked on the clean CIFAR10. We report the mean accuracy over 30 runs. LA-SiLU achieved the best performance among the activation functions. In other experiments, with ResNet32 and ResNet44 on clean CIFAR10 (Appendix J), networks with GELU achieved better results in two cases: ResNet32 on CIFAR10 and ResNet44 on CIFAR100. In the remaining combinations of networks and datasets, LA-SiLU outperformed in a significant majority of cases, especially in 39 out of 48 cases according to the $p$-value from a T-test or a Wilcoxon signed-rank test, indicating statistical significance.
@@
Table 3 shows the performance of ResNet50 and ResNet101 with LayerAct functions and other ElementAct functions for comparison with clean and noisy ImageNet datasets. For the column named ”clean,” we report the accuracy of 10-crop testing on the validation dataset. We used the out-of-distribution benchmark dataset, ImageNet-C (Hendrycks and Dietterich 2019), which has the same corruptions as CIFAR-C datasets. The networks with LayerAct functions outperformed those with other activation functions on all datasets. The ResNet50 with LayerAct functions, even with LA-HardSiLU which showed worse performance on the clean CIFAR10 datasets compared to SiLU or LReLU, outperformed other activation functions on clean ImageNet. Based on these results, we trained deeper networks, ResNet101, utilizing ReLU and LA-SiLU as activations. LA-SiLU maintained its superior robustness.